[Zach and Kiwamu]
We worked on some more vacuum businesses. Today we finished did the following works:
- alignment of the POX mirrors
- alignment of the POP1 and POP2 mirrors
- installation of OSEMs onto SRM
- alignment of the SRM tower
(alignment of POP mirrors)
Since a beam on the POP path was quite too weak to see even by IR viewers, we used a He-Ne laser to imitate the real beam instead.
We injected the He-Ne beam from an optical bench to the chamber, and made it go through the PRM and PR2 by using some steering mirrors.
The pin assignment was flipped in a way of mirror image due to the extension cables which cause a mirroring.
So we made mirroring connectors to flipp them back to the correct pin assignment, and plugged the mirroring connectors in between the feedthrough of the BS chamber and the SRM satellite box.
This is a picture showing how they are connected now.